What is the solution? The square root of 1 - y = 8
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

sqrt(1-y)=8
square both sides
1-y=64
y=-63..

Answer by stanbon(75887)

What is the solution? The square root of 1 - y = 8
----------------
sqrt(1-y) = 8
---
Square both sides to get:
1-y = 64
y = 1-64
y = -63
------------------
Check that "answer".
sqrt(1--63)
= sqrt(64)
= 8
======
Final answer: y = -63
============================
Cheers,
Stan H.